Errata in my Diploma Thesis A Semantical Analysis of Structural Recursion

Andreas Abel October 12, 1999

p. 17 Operational semantics There is an error in the rule (opappvr). The correct rule is

(opappvr) {⟨t; e, g = ⟨rec g.t; e′⟩⟩ ⇓ f f@u ⇓ v \atop ⟨rec g.t; e⟩@u ⇓ v}